# Name That Part: 3D Part Segmentation and Naming

### v1

@@
Abstract.
We address semantic 3D part segmentation: decomposing objects into parts with meaningful names. While datasets exist with part annotations, their definitions are inconsistent across datasets, limiting robust training. Previous methods produce unlabeled decompositions or retrieve single parts without complete shape annotations. We propose ALIGN-Parts, which formulates part naming as a direct set alignment task. Our method decomposes shapes into partlets - implicit 3D part representations - matched to part descriptions via bipartite assignment. We combine geometric cues from 3D part fields, appearance from multi-view vision features, and semantic knowledge from language-model-generated affordance descriptions. Text-alignment loss ensures partlets share embedding space with text, enabling a theoretically open-vocabulary matching setup, given sufficient data. Our efficient and novel, one-shot, 3D part segmentation and naming method finds applications in several downstream tasks, including serving as a scalable annotation engine. As our model supports zero-shot matching to arbitrary descriptions and confidence-calibrated predictions for known categories, with human verification, we create a unified ontology that aligns PartNet, 3DCoMPaT++, and Find3D, consisting of 1,794 unique 3D parts. We also show examples from our newly created Tex-Parts dataset. We also introduce 2 novel metrics appropriate for the named 3D part segmentation task.
3D part semantic segmentation, geometric deep learning
copyright: none
@@ Abstract.
Table 1. Comparison with 3D part methods.
1. Introduction
Many vision tasks require 3D parts, not just whole-object labels. Examples include robots grasping handles and creators editing assets. This requires solving two problems simultaneously: geometrically segmenting the parts and semantically naming them. While large datasets of 3D objects exist, only a few provide part annotations, and these annotations are often inconsistent across datasets (Mo_2019_CVPR; slim_3dcompatplus_2023; ma2024find). An algorithm that can provide accurate and consistent annotations of named parts on any 3D object would enable scalable training data and support human-in-the-loop annotation pipelines.
Existing methods address only one aspect of this problem. Part segmentation models can identify geometric boundaries but produce unnamed regions (liu2025partfield). Language-grounded systems can retrieve a single part from a text query but fail to produce a complete set of names for all parts of an object (ma2024find). Classical unsupervised part discovery lacks the semantic grounding to derive consistent definitions. This gap has created a bottleneck: the absence of large-scale, consistently-annotated 3D part data. Inducing consistent parts from unlabeled web assets requires coupling geometric features and semantic knowledge with human verification.
Figure 2. ALIGN-Parts segments and names 3D parts, unlike PartField (liu2025partfield), which only segments. Our method is $100\times$ (post data pre-processing) faster at generating these segments along with names compared to PartField.
@@ 1. Introduction
Our key contribution is applying set-level matching to 3D part fields. This enables lower computational complexity: we match a handful of partlets to descriptions instead of all points to all labels. It provides permutation consistency: each predicted part receives at most one name, and each name is used at most once. Because of this, ALIGN-Parts is 100× faster at generating these segments along with names.
To ensure partlets are both geometrically separable and semantically meaningful, we combine three signals. (1) Geometry features from a 3D part-field backbone (liu2025partfield) capture shape structure. (2) Appearance features from multi-view image encoders (oquab2023dinov2) lifted onto 3D geometry provide texture cues.
(3) Semantic knowledge comes from affordance-aware part descriptions that encode form-and-function relationships (connell1987generating). For example, a chair seat becomes “the horizontal surface where a person sits,” linking its flat, horizontal geometry to its sitting affordance.
For text grounding, we generate these descriptions using a large language model (comanici2025gemini) and embed them using MPNet sentence transformers (song2020mpnet). By representing part descriptions as embeddings in a continuous space, our approach supports scenarios where the model can match partlets to any user-provided set of part descriptions without retraining.
Our approach provides a tool to address the data bottleneck. We construct a unified part ontology using a hybrid LLM-and-human process that normalizes labels and verifies geometric consistency across PartNet, 3DCoMPaT++, and Find3D. We then apply ALIGN-Parts to bootstrap annotations from unlabeled TexVerse assets, creating Tex-Parts: a dataset with 8450 objects spanning 14k part categories. ALIGN-Parts serves as a scalable annotation engine that proposes named parts for human verification, converting raw meshes into training data with minimal effort and enabling the construction of large, consistently-annotated 3D datasets.
@@
  • Unified ontology and scalable annotation engine. We harmonize part taxonomies across PartNet, 3DCoMPaT++, and Find3D, and demonstrate a human-in-the-loop pipeline that bootstraps TexParts, a verified benchmark of 8450 objects spanning 14k categories derived from Texverse (zhang2025texverse). ALIGN-Parts converts raw meshes into training data with $5$-$8\times$ less human effort than manual annotation.
2. Related Work
@@
3D Part Segmentation. Traditional methods operate in a purely geometric regime. PartField (liu2025partfield), the current state-of-the-art, learns dense per-point feature fields but produces unlabeled regions. Recent works lift 2D foundation models into 3D: SAM-based approaches (ma2025p3sam; yang2024sampart3d; tang2024segment) adapt Segment-Anything via multi-view projection but require prompts and lack semantic names. PartSTAD (kim2024partstad) adapts GLIP and SAM with 3D-aware objectives, while Diff3F (dutt2024diffusion) exploits diffusion features, but both operate as multi-stage pipelines without producing complete, non-overlapping named parts.
To our knowledge, (Kalogerakis:2010:labelMeshes; Kalogerakis2017) are the only previous works that do simultaneous 3D part segmentation and labeling. (Kalogerakis:2010:labelMeshes) defines the task by assigning fixed labels to mesh faces using statistical consistency using a Conditional Random Field Model, which is hard to scale up. Further, the model lacks any semantic knowledge given its the reliance on hand engineering features and geometric consistency. Simultaneous 3D part segmentation and recognition has remained a complex task with little progress since then as it requires incorporation of human contextual knowledge of parts. Therefore, most subsequent works have focused on amodal segmentation, which is a relatively easier task.
Other important works in the area of 3D parts - which are not directly relevant to our work - but involve relevant 3D analysis, understanding or retrieval, include (10.1007/s00371-007-0197-5; jones2020shapeAssembly; 10.1145/1409060.1409098; hanocka2019meshcnn; abstractionTulsiani17; Stelzner2021Decomposing3S; takmaz2023openmask3d).
@@
Language-Grounded 3D Understanding. PartSLIP (liu2023partslip), PartSLIP++ (zhou2023partslip++), and PartDistill (umam2023partdistill) use image-language models for part segmentation but require per-category fine-tuning with predefined vocabularies. Find3D (ma2024find) is most related: it trains a point transformer in text embedding space for text-to-part retrieval. However, Find3D operates query-by-query (given "wing, head," it returns masks) rather than producing complete decompositions, and works on individual point features rather than part-level aggregations.
@@
ALIGN-Parts differs fundamentally. We formulate part segmentation and naming as direct set alignment: Partlets aggregate point features into part-level representations matched to text via optimal transport, enabling (1) complete, non-overlapping decompositions in one pass, (2) permutation-consistent naming, (3) dynamic part counts, and (4) zero-shot generalization (Table 1). Unlike previous methods, ALIGN-Parts jointly learns segmentation and alignment in a single feed-forward pass. Part count emerges automatically from activated Partlets, eliminating dependence on K prediction or part-name inputs.
@@
Part-Based Generation and Datasets. Generative methods (chen2024partgen; chen2025autopartgen; tang2024partpacker; lin2025partcrafter; yang2025holopart) discover part structure through synthesis but lack open-vocabulary naming mechanisms. For datasets, PartNet (Mo_2019_CVPR), 3DCoMPaT++ (slim_3dcompatplus_2023), and GAPartNet (geng2023gapartnet) provide part annotations but use inconsistent taxonomies. ALIGN-Parts constructs a unified ontology across these datasets and creates a unified dataset for cross-dataset evaluation.
3. Method
@@
Figure 3. ALIGN-Parts. Overview of the ALIGN-Parts framework for language-grounded 3D part segmentation and naming. Top: training. Given a 3D from our semantically unified 3D parts data, geometry features are extracted with PartField and appearance features with DINOv2 from multi-view renderings; these are fused by the BiCo Fusion module using efficient bi-directional cross-attention on local k=16k{=}16 nearest-neighbor graphs in 3D space, reducing complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(Nk)$ and yielding enriched point features. A decoder then learns $K$ part-level “Partlet” representations that aggregate the fused features, with segmentation supervision provided at the Partlet level. To semantically ground Partlets, an LLM generates affordance-aware descriptions for each possible part (e.g., “structural supports that elevate the sofa” for sofa legs), which are embedded by a pretrained MPNet encoder; Sinkhorn matching establishes a bipartite assignment between Partlet and text embeddings, and an InfoNCE loss further aligns them in a shared representation space while a classifier predicts the object category. Bottom: inference. At test time, ALIGN-Parts operates in both closed-vocabulary (object categories similar to those seen in training) and open-vocabulary (novel object categories) settings: in the closed-vocabulary case, the trained 3D classifier predicts the object class and retrieves its candidate part list, whereas in the open-vocabulary case an LLM proposes an overcomplete set of plausible parts for the queried object. Given these candidate part texts, their MPNet embeddings are bipartite-matched to the predicted Partlets, which jointly produce 3D part segmentation masks and corresponding part names.
@@
Overview. We propose ALIGN-Parts, a framework that treats automatic semantic 3D part segmentation as a direct set alignment problem, analogous to DETR (carion2020end) in 2D detection. The pipeline consists of three components: (1) a dense feature fusion module combining geometry and appearance via localized bi-directional 3D-aware cross-attention, (2) a Partlets module that learns K adaptive part-level representations by aggregating fused point features, and (3) a semantic grounding module that aligns Partlets to part names via text descriptions of part affordances. At inference, given a 3D shape and candidate part descriptions, our method produces named part segments without specifying part counts, names, or point prompts. Training uses contrastive alignment and differentiable optimal transport matching between predicted Partlets and ground-truth parts, enabling end-to-end learning of both segmentation and semantics.
@@
Problem Formulation. Given a 3D point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}$, each point has geometry-first features $\mathbf{f}_{i}^{g}\in\mathbb{R}^{d_{g}}$, appearance-first features $\mathbf{f}_{i}^{a}\in\mathbb{R}^{d_{a}}$, and coordinates $\mathbf{x}_{i}\in\mathbb{R}^{3}$. During training, we have ground-truth part masks $\mathbf{M}^{\text{gt}}\in\{0,1\}^{A\times N}$ and corresponding text embeddings $\{\mathbf{t}_{a}\}_{a=1}^{A}$ where $\mathbf{t}_{a}\in\mathbb{R}^{d_{t}}$ is an MPNet embedding of a language-model-generated part description.
@@
Our model learns $K=32$ instance-specific Partlet embeddings $\{\mathbf{s}_{k}\in\mathbb{R}^{d_{t}}\}_{k=1}^{K}$ that reside in the same semantic space as text embeddings (we use $d_{t}=768$). Each Partlet predicts three outputs: (i) a soft mask $\mathbf{m}_{k}\in[0,1]^{N}$ over points, (ii) a partness score $\text{part}_{k}\in\mathbb{R}$ indicating whether it represents an actual part, and (iii) its embedding $\mathbf{s}_{k}$ which serves as the prototype for semantic alignment. We select $K=32$ to accommodate variable part configurations (most shapes have $\leq 28$ parts) while allowing future extensibility. Point labels are obtained by matching Partlets to ground-truth parts via optimal transport.
3.1. Architecture
@@
Dense Feature Fusion Module The raw geometric features $\mathbf{f}_{i}^{g}$ and appearance features $\mathbf{f}_{i}^{a}$ capture complementary information: geometry encodes shape characteristics, while appearance provides texture and visual cues. We fuse these modalities through a bi-directional cross-attention module that operates on a $k$-nearest neighbor (KNN) graph to maintain computational tractability.
Relative Positional Bias.
To incorporate 3D spatial structure through Fourier-encoded relative positional biases, for each neighbor pair $(i,j)$ where $j\in\mathcal{N}_{i}$, we compute:
(1) $\mathbf{d}_{ij}$ $=\mathbf{x}_{j}-\mathbf{x}_{i}\in\mathbb{R}^{3}$
(2) $\mathbf{f}_{ij}$ $=\mathbf{d}_{ij}\odot\omega\in\mathbb{R}^{3\times F}$
(3) $\mathbf{h}_{ij}$ $=[\sin(\mathbf{f}_{ij}),\cos(\mathbf{f}_{ij})]\in\mathbb{R}^{3\times 2F}$
(4) $\mathbf{b}_{ij}$ $=\text{MLP}([\mathbf{d}_{ij},\text{flatten}(\mathbf{h}_{ij})])\in\mathbb{R}^{H}$
Bi-Directional Cross-Attention.
Geometric-to-Appearance Direction:
(5) $\mathbf{Q}_{p}^{i}$ $=\mathbf{W}_{q}^{p}\mathbf{f}_{i}^{g}\in\mathbb{R}^{H\times d_{h}}$
(6) $\mathbf{K}_{a}^{ij}$ $=\mathbf{W}_{k}^{a}\mathbf{f}_{j}^{a}\in\mathbb{R}^{H\times d_{h}},\quad\forall j\in\mathcal{N}_{i}$
(7) $\mathbf{V}_{a}^{ij}$ $=\mathbf{W}_{v}^{a}\mathbf{f}_{j}^{a}\in\mathbb{R}^{H\times d_{h}},\quad\forall j\in\mathcal{N}_{i}$
(8) $\alpha_{pa}^{i,h,j}$ $=\frac{\exp((\mathbf{Q}_{p}^{i}[h]\cdot\mathbf{K}_{a}^{ij}[h])/\sqrt{d_{h}}+\mathbf{b}_{ij}[h])}{\sum_{j^{\prime}\in\mathcal{N}_{i}}\exp((\mathbf{Q}_{p}^{i}[h]\cdot\mathbf{K}_{a}^{ij^{\prime}}[h])/\sqrt{d_{h}}+\mathbf{b}_{ij^{\prime}}[h])}$
(9) $\mathbf{z}_{p}^{i,h}$ $=\sum_{j\in\mathcal{N}_{i}}\alpha_{pa}^{i,h,j}\mathbf{V}_{a}^{ij}[h]$
(10) $\mathbf{r}_{p}^{i}$ $=\mathbf{W}_{pa}\text{concat}_{h}[\mathbf{z}_{p}^{i,h}]\in\mathbb{R}^{d_{g}}$
Appearance-to-Geometric Direction:
Gated Fusion.
(11) $\mathbf{g}_{p}^{i}$ $=\sigma(\mathbf{W}_{g}^{p}[\mathbf{f}_{i}^{g};\mathbf{r}_{p}^{i}])\in\mathbb{R}^{d_{g}}$
(12) $\mathbf{g}_{a}^{i}$ $=\sigma(\mathbf{W}_{g}^{a}[\mathbf{f}_{i}^{a};\mathbf{r}_{a}^{i}])\in\mathbb{R}^{d_{a}}$
(13) $\tilde{\mathbf{f}}_{i}^{g}$ $=\text{LayerNorm}(\mathbf{f}_{i}^{g}+\mathbf{g}_{p}^{i}\odot\mathbf{r}_{p}^{i})$
(14) $\tilde{\mathbf{f}}_{i}^{a}$ $=\text{LayerNorm}(\mathbf{f}_{i}^{a}+\mathbf{g}_{a}^{i}\odot\mathbf{r}_{a}^{i})$
Final Projection.
(15) $\mathbf{h}_{i}=\mathbf{W}_{2}\text{GELU}(\mathbf{W}_{1}\text{LayerNorm}([\tilde{\mathbf{f}}_{i}^{g};\tilde{\mathbf{f}}_{i}^{a}]))$
@@
Points $\rightarrow$ Partlets: Learning Part-Level Representations. We learn K Partlet embeddings that aggregate point-level information into part-level representations. A Partlet is defined by three components: (i) a soft segmentation mask $\mathbf{m}\in[0,1]^{N}$ representing membership scores, (ii) a Partlet embedding $\mathbf{s}\in\mathbb{R}^{d_{t}}$ in the learned semantic space, and (iii) a text embedding $\mathbf{z}\in\mathbb{R}^{d_{t}}$ representing the part description.
@@
(16) $(\mathbf{m}_{k},\mathbf{s}_{k})=f_{\theta}(\mathbf{H}),\quad k=1,\ldots,K$
@@
(17) $\mathbf{s}_{k}^{(\ell,1)}=\mathbf{s}_{k}^{(\ell-1)}+\text{SelfAttn}(\mathbf{s}_{k}^{(\ell-1)},\{\mathbf{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K})$
@@
(18) $\mathbf{s}_{k}^{(\ell,2)}=\mathbf{s}_{k}^{(\ell,1)}+\text{CrossAttn}(\mathbf{s}_{k}^{(\ell,1)},\{\mathbf{h}_{i}\}_{i=1}^{N})$
@@
(19) $\mathbf{s}_{k}^{(\ell)}=\mathbf{s}_{k}^{(\ell,2)}+\text{MLP}(\mathbf{s}_{k}^{(\ell,2)})$
@@
Mask Prediction. Each Partlet predicts point membership via scaled dot-product:
(20) $m_{ki}=\frac{(\mathbf{W}_{q}\mathbf{s}_{k})^{T}(\mathbf{W}_{k}\mathbf{h}_{i})}{\sqrt{d_{t}}}$
@@
Partness Prediction. Each Partlet predicts whether it represents an actual part:
(21) $\text{part}_{k}=\mathbf{w}_{\text{part}}^{T}\mathbf{s}_{k}+b_{\text{part}}$

### v2

@@
Abstract
We address semantic 3D part segmentation: decomposing objects into parts with meaningful names. While datasets exist with part annotations, their definitions are inconsistent across datasets, limiting robust training. Previous methods produce unlabeled decompositions or retrieve single parts without complete shape annotations. We propose ALIGN-Parts, which formulates part naming as a direct set alignment task. Our method decomposes shapes into partlets - implicit 3D part representations - matched to part descriptions via bipartite assignment. We combine geometric cues from 3D part fields, appearance cues from multi-view vision features, and semantic knowledge from language-model-generated affordance descriptions. Text-alignment loss ensures partlets share embedding space with text, enabling a theoretically open-vocabulary matching setup, given sufficient data. Our efficient and novel, one-shot, 3D part segmentation and naming method finds applications in several downstream tasks, including serving as a scalable annotation engine. As our model supports zero-shot matching to arbitrary descriptions and confidence-calibrated predictions for known categories, with human verification, we create a unified ontology that aligns PartNet, 3DCoMPaT++, and Find3D, consisting of 1,794 unique 3D parts. We introduce two novel metrics appropriate for the named 3D part segmentation task. We also show examples from our newly created TexParts dataset.
Figure 1: ALIGN-Parts is the first large-scale method to be able to efficiently and semantically segment and name 3D parts of an object, unlike previous methods, which could only perform class-agnostic part segmentation. Our method is also feed-forward, and defines 3D parts according to human-oriented, object-specific affordance descriptions. (left) ALIGN-Parts segments and semantically names 3D parts, unlike PartField [23], which only segments. Our feed-forward method is faster at generating these segments along with names compared to PartField as we don’t require clustering. (right) Simultaneous part segmentation and naming enable $5-8\times$ more efficient creation of 3D part datasets (bottom) with human-in-the-loop validation.
@@ Abstract
Table 1: Comparison with 3D Part segmentation or retrieval methods.
1 Introduction
Many vision tasks require 3D parts, not just whole-object labels. Examples include robots grasping handles and creators editing assets. This requires solving two problems simultaneously: geometrically segmenting the parts and semantically naming them. While large datasets of 3D objects exist, only a few provide part annotations, and these annotations are often inconsistent across datasets [27, 31, 26]. An algorithm that can provide accurate and consistent annotations of named parts on any 3D object would enable scalable training data and support human-in-the-loop annotation pipelines.
Existing methods address only one aspect of this problem. Part segmentation models can identify geometric boundaries but produce unnamed regions [23]; however, these arbitrary segmentations lack any semantic grounding (from a part-based human perspective). Language-grounded systems can retrieve a single part from a text query but fail to produce a complete set of names for all parts of an object [26]. Classical unsupervised part discovery lacks the semantic grounding to derive consistent definitions. This gap has created a bottleneck: the absence of large-scale, consistently-annotated 3D part data. Inducing consistent parts from unlabeled web assets requires coupling geometric features and semantic knowledge with human verification.
@@ 1 Introduction
Our key contribution is the application of set-level matching to 3D part fields. This enables lower computational complexity: we match a handful of partlets to descriptions instead of all points to all labels. It provides permutation consistency: each predicted part receives at most one name, and each name is used at most once. Because of this, ALIGN-Parts is significantly faster at generating these segments, along with their names.
To ensure partlets are both geometrically separable and semantically meaningful, we combine three signals -
  1. 1.
    Geometry features from a 3D part-field backbone [23] capture shape structure,
  2. 2.
    Appearance features from multi-view image encoders [28] lifted onto 3D geometry provide texture cues,
  3. 3.
    Semantic knowledge comes from affordance-aware part descriptions that encode form-and-function relationships [6]. For example, a chair seat becomes “the horizontal surface where a person sits,” linking its flat, horizontal geometry to its sitting affordance.
For semantic text grounding, we generate affordance and category-based part descriptions using a large language model [5] and embed them using MPNet sentence transformers [32]. By representing part descriptions as embeddings in a continuous space, our approach supports scenarios where the model can match partlets to any user-provided set of part descriptions without retraining. In order to evaluate the task of simultaneous 3D part segmentation and naming, we introduce two new metrics.
Our approach provides a tool to address the 3D part data bottleneck. We construct a unified part ontology using a hybrid LLM-and-human process that normalizes labels and verifies geometric consistency across PartNet, 3DCoMPaT++, and Find3D. We then apply ALIGN-Parts to bootstrap annotations from unlabeled TexVerse assets [43], creating Tex-Parts: a dataset with 8450 objects spanning 14k part categories. In this setup, ALIGN-Parts serves as a scalable annotation engine that proposes named parts for human verification, converting raw meshes into training data with minimal effort and enabling the construction of large, consistently-annotated 3D datasets.
@@
    We harmonize part taxonomies across PartNet, 3DCoMPaT++, and Find3D datasets, and demonstrate a human-in-the-loop pipeline that bootstraps TexParts, a verified benchmark of 8450 objects spanning 14k categories derived from Texverse [43]. ALIGN-Parts converts raw meshes into training data with $5$-$8\times$ less human effort than manual annotation.
2 Related Work
3D Part Segmentation.
Traditional methods operate in a purely geometric regime [30, 10]. Early methods rely on handcrafted features and geometric consistency to partition shapes into meaningful regions, without assigning semantic names. Recent advances, such as PartField [23], the current state-of-the-art, learns dense per-point feature fields but produce unlabeled regions.
@@ 2 Related Work
More recent works lift 2D foundation models into 3D: SAM-based approaches [25, 42, 35] adapt Segment-Anything via multi-view projection but require prompts and lack semantic names. PartSTAD [17] integrates GLIP and SAM with 3D-aware objectives, while Diff3F [8] exploits diffusion features for segmentation. However, these methods operate as multi-stage pipelines and do not produce complete, non-overlapping part decompositions with coherent semantic 3D part names.
Among prior works, Kalogerakis et al. [15, 14] are the closest to our problem setting, as they explicitly formulate joint 3D part segmentation and semantic labeling. [15] assigns fixed semantic labels to mesh faces using a Conditional Random Field with handcrafted geometric features, while the [14] extension introduces learned features via projective convolutions. Despite their importance, these approaches rely on closed, category-specific label sets and lack the ability to scale to open-vocabulary naming or instance-consistent part identities. Moreover, labeling is framed as classification over predefined semantics rather than as a language-grounded naming problem.
As a result, most subsequent work has focused on amodal or class-based part segmentation, avoiding the challenges of assigning coherent and permutation-consistent semantic names to discovered parts. Other works related to 3D part analysis, but not directly addressing part naming, include shape correspondence, retrieval, and structural understanding methods [13, 12, 37, 33, 34].
Language-Grounded 3D Understanding.
@@
PartSLIP [24], PartSLIP++ [44], and PartDistill [39] use image-language models for part segmentation but require per-category fine-tuning with predefined vocabularies. Find3D [26] and PartGlot [19] are most related: Find3D trains a point transformer in text embedding space for text-to-part retrieval. However, Find3D operates query-by-query (given "wing, head," it returns masks) rather than producing complete decompositions, and works on individual point features rather than part-level aggregations. PartGlot derives part-level segmentation masks from spatial attention as a byproduct of learning to play the language reference game. However, it is formulated as a choice game for only a single class of object (chair), and a predefined small number of parts.
ALIGN-Parts
@@
ALIGN-Parts differs fundamentally from prior work by formulating part segmentation and naming as a set alignment problem: Partlets aggregate point features into part-level representations matched to text via optimal transport. This formulation enables: (1) complete, non-overlapping part decompositions in a single forward pass, (2) permutation-consistent semantic naming, (3) dynamic part cardinality without predefined part-count, and (4) zero-shot generalization across categories and datasets (Tab.˜1). Unlike prior multi-stage or query-based methods, ALIGN-Parts jointly learns segmentation and semantic alignment in an end-to-end manner. The number of parts emerges automatically from activated Partlets, eliminating the need for explicit part-count prediction or part-name inputs at inference time.
Part-Based Generation and Datasets.
@@
Part-based generative models [3, 4, 36, 22, 41] discover latent part structures through synthesis, but do not address open-vocabulary semantic naming. Existing datasets such as PartNet [27], 3DCoMPaT++ [21, 31], and GAPartNet [9] provide part annotations, but rely on inconsistent taxonomies and category-specific semantics. ALIGN-Parts constructs a unified ontology across these datasets, enabling cross-dataset evaluation under a shared semantic framework.
@@
Figure 2: ALIGN-Parts. Overview of the ALIGN-Parts framework for language-grounded 3D part segmentation and naming. Top: training. Given a 3D from our semantically unified 3D parts data, geometry features are extracted with PartField and appearance features with DINOv2 from multi-view renderings; these are fused by the BiCo Fusion module using efficient bi-directional cross-attention on local k=16k{=}16 nearest-neighbor graphs in 3D space, reducing complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(Nk)$ and yielding enriched point features. A decoder then learns $K$ part-level “Partlet” representations that aggregate the fused features, with segmentation supervision provided at the Partlet level. To semantically ground Partlets, an LLM generates affordance-aware descriptions for each possible part (e.g., “structural supports that elevate the sofa” for sofa legs), which are embedded by a pretrained MPNet encoder; Sinkhorn matching establishes a bipartite assignment between Partlet and text embeddings, and an InfoNCE loss further aligns them in a shared representation space while a classifier predicts the object category. Bottom: inference. At test time, ALIGN-Parts operates in both closed-vocabulary (object categories similar to those seen in training) and open-vocabulary (novel object categories) settings: in the closed-vocabulary case, the trained 3D classifier predicts the object class and retrieves its candidate part list, whereas in the open-vocabulary case an LLM proposes an overcomplete set of plausible parts for the queried object. Given these candidate part texts, their MPNet embeddings are bipartite-matched to the predicted Partlets, which jointly produce 3D part segmentation masks and corresponding part names.
3 Method
Overview.
We propose ALIGN-Parts, a framework that treats automatic semantic 3D part segmentation as a direct set alignment problem, analogous to DETR [2] in 2D detection. The pipeline consists of three components:
  1. 1.
    A dense feature fusion module combining geometry and appearance via localized bi-directional 3D-aware cross-attention,
  2. 2.
    A Partlets module that learns K adaptive part-level representations by aggregating fused point features, and
  3. 3.
    A semantic grounding module that aligns Partlets to part names via text descriptions of part affordances.
@@ 3 Method
At inference time, given a 3D shape and generated candidate part descriptions, our method produces named part segments without specifying part counts, names, or point prompts. Training uses contrastive alignment and differentiable optimal transport matching between predicted Partlets and ground-truth parts, enabling end-to-end learning of both segmentation and semantics.
Problem Formulation.
@@
Given a 3D point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}$, each point has geometry-first features $\mathbf{f}_{i}^{g}\in\mathbb{R}^{d_{g}}$, appearance-first features $\mathbf{f}_{i}^{a}\in\mathbb{R}^{d_{a}}$, and coordinates $\mathbf{x}_{i}\in\mathbb{R}^{3}$. During training, we have ground-truth part masks $\mathbf{M}^{\text{gt}}\in\{0,1\}^{A\times N}$ and corresponding text embeddings $\{\mathbf{t}_{a}\}_{a=1}^{A}$ where $\mathbf{t}_{a}\in\mathbb{R}^{d_{t}}$ is an MPNet embedding of a language-model-generated semantic part description.
@@
Our model learns $K=32$ instance-specific Partlet embeddings $\{\mathbf{s}_{k}\in\mathbb{R}^{d_{t}}\}_{k=1}^{K}$ that reside in the same dimensional space as semantic text embeddings (we use $d_{t}=768$). Each Partlet predicts three outputs: (i) a soft part mask $\mathbf{m}_{k}\in[0,1]^{N}$ over points, (ii) a partness score $\text{part}_{k}\in\mathbb{R}$ indicating whether it represents an actual part, and (iii) its embedding $\mathbf{s}_{k}$ which serves as the prototype for semantic alignment. We select $K=32$ to accommodate variable part configurations (We find that most 3D shapes have $\leq 28$ parts) while allowing future extensibility. Point part labels are obtained by matching Partlets to ground-truth or generated parts’ descriptions via optimal transport.
3.1 Architecture
The input to our model is raw geometric and appearance features. We use PartField and multiview DiNO features; however, they can be replaced with any viable alternatives. Ideally, we want to train these input feature modules in an end-to-end manner with the rest of our model to extract optimal performance from the model. However, we are constrained due to compute requirements and lack of sufficient finetuning details [23]. Better performance than shown in this work is likely possible with better and trainable input modules.
Feature Fusion.
@@ 3.1 Architecture
The raw geometric features $\mathbf{f}_{i}^{g}$ and appearance features $\mathbf{f}_{i}^{a}$ capture complementary information: geometry encodes shape characteristics, while appearance provides texture and visual cues. We fuse these modalities through bi-directional cross-attention that we call BiCo Fusion operating on local k=16k{=}16 nearest neighbor graphs in 3D coordinate space, reducing complexity to $\mathcal{O}(Nk)$.
For each point $i$ along with its nearest neighbors, geometric features attend to the appearance features of neighbors, producing cross-modal features that capture appearance information. Symmetrically, appearance features attend to geometric features of neighbors, capturing geometric information. We incorporate 3D spatial structure through Fourier-encoded relative positional biases. Learned sigmoid gates control how much of this cross-modal information to incorporate into each original feature, based on both the original feature and the attended information. After gated addition and layer normalization, we concatenate both modalities and project through a two-layer MLP to produce fused features $\mathbf{h}_{i}\in\mathbb{R}^{256}$ for each point. See Sec.˜A.1 for details.
Points $\rightarrow$ Partlets: Learning Part-Level Representations.
@@
We learn $K$ Partlet embeddings that aggregate point-level information into part-level representations. A Partlet is defined by three components: (i) a soft segmentation mask $\mathbf{m}\in[0,1]^{N}$ representing membership scores, (ii) a Partlet embedding $\mathbf{s}\in\mathbb{R}^{d_{t}}$ in the learned semantic space, and (iii) a text embedding $\mathbf{z}\in\mathbb{R}^{d_{t}}$ representing the part description.
@@
$(\mathbf{m}_{k},\mathbf{s}_{k})=f_{\theta}(\mathbf{H}),\quad k=1,\ldots,K$ (1)
@@
$\mathbf{s}_{k}^{(\ell,1)}=\mathbf{s}_{k}^{(\ell-1)}+\text{SelfAttn}(\mathbf{s}_{k}^{(\ell-1)},\{\mathbf{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K})$ (2)
@@
$\mathbf{s}_{k}^{(\ell,2)}=\mathbf{s}_{k}^{(\ell,1)}+\text{CrossAttn}(\mathbf{s}_{k}^{(\ell,1)},\{\mathbf{h}_{i}\}_{i=1}^{N})$ (3)
@@
$\mathbf{s}_{k}^{(\ell)}=\mathbf{s}_{k}^{(\ell,2)}+\text{MLP}(\mathbf{s}_{k}^{(\ell,2)})$ (4)
@@
Mask Prediction.
Each Partlet predicts point membership via scaled dot-product:
$m_{ki}=\frac{(\mathbf{W}_{q}\mathbf{s}_{k})^{T}(\mathbf{W}_{k}\mathbf{h}_{i})}{\sqrt{d_{t}}}$ (5)
@@
Partness Prediction.
Each Partlet predicts whether it represents an actual part:
$\text{part}_{k}=\mathbf{w}_{\text{part}}^{T}\mathbf{s}_{k}+b_{\text{part}}$ (6)